# Adversarial Distilled Retrieval-Augmented Guarding Model for Online Malicious Intent Detection

### v1

@@ 5.1 RQ1 – Does ADRAG yield more accurate and faster models?
Metric / Dataset
BERT
SFT (395M)
GPT-4
WildGuard
Llama-Guard-3 (8B)
ADRAG
ADRAG
SKD (149M)
@@
Table 1: F1 scores and latency of ADRAG vs. baselines. Best results in bold; GPT-4 latency not reported.
On ID datasets, ADRAG surpasses all baselines on Dynahate (0.889 vs. 0.752 for Llama-Guard-3) and remains competitive on WildMix (0.893 vs. 0.905), even though WildGuard was trained on WildMix. On OOD datasets, ADRAG generally matches or exceeds large baselines. It performs competitively on ToxicChat (0.921 vs. 0.925/0.932) and ToxiGen (0.821 vs. 0.743/0.829), and clearly outperforms on JailJudge (0.650 vs. 0.436/0.531). Performance is close but slightly lower on WildJailbreak (0.932 vs. 0.964) and CounterfactualFair (0.889 vs. 0.907/0.935), where ultra-large models retain a modest edge.
The distilled variant, ADRAG-SKD (149M), preserves most of this accuracy while further reducing model size. For efficiency, ADRAG achieves $p_{90}$$p_{99}$ latencies of 4.38/4.43 ms, dramatically faster than WildGuard (46.70/50.55 ms) and Llama-Guard-3 (43.82/50.70 ms). ADRAG-SKD reduces latency further to 2.83/3.07 ms with minimal accuracy loss. In online deployment, retrieval adds only $\sim$4 ms, keeping end-to-end latency under 9 ms (see Appendix A.9). These results demonstrate that ADRAG delivers large-model accuracy at a fraction of the size and latency, while maintaining robustness across both ID and OOD settings.
@@ 5.2 RQ2 – Contribution of RAFT and SKD
We compare four configurations under a more controllable setting, trained only on WildJailbreak: a supervised baseline (SFT), Retrieval-Augment Fine-Tuning (RFT), Retrieval-Augmented Adversarial Fine-Tuning (RAFT), and Selective Knowledge Distillation (SKD). The experiments show that RAG enhances generalization, RAFT strengthens robustness and accuracy, and SKD offers the best accuracy–efficiency trade-off.
As shown in Table 2, RFT improves on Prosocial (+0.031) and modestly on WildJailbreak, WildMix, and ToxiGen, but regresses on Dynahate (0.057-0.057), revealing vulnerability to adversarial contexts 3. RAFT boosts robustness and achieves the best overall gain (+0.043), with notable gains on ToxiGen (+0.056) and Prosocial (+0.071). Even on Dynahate, it reduces the drop to 0.020-0.020, showing better tolerance to noisy supervision. SKD, despite its smaller size, achieves a similar overall gain (+0.035) and surpasses RAFT on ToxicChat (+0.065), proving that distillation preserves key knowledge while lowering inference latency.
F1-score
RAFT / SKD
$\Delta$ vs. SFT
Dataset
SFT
395M
RFT
395M
RAFT
395M
SKD
149M
@@ 5.3 RQ3 – Further Discussion
Our ablation studies yield several additional insights. Appendix A.3 shows that tuning the retrieval component alone is inadequate, underscoring the need for robust context-aware modeling through targeted training and motivating the introduction of RAFT. Appendix A.2 reveals that smaller models often struggle with context-dependent patterns and domain generalization, while SKD helps narrow this gap. Appendix A.4 provides guidance on tuning SKD loss functions, particularly the trade-off between Kullback–Leibler (KL) divergence and cross-entropy. Appendix A.5 demonstrates that ADRAG consistently improves performance across model sizes and architectures, with stronger base model (e.g., ModernBERT versus Google BERT) benefiting the most. Finally, large-scale hyperparameter exploration (over 2k trials) indicates that ADRAG offers limited gains in domains where the baseline already approaches ceiling performance (e.g., F1 $\geq$ 0.95 on WildJailbreak), suggesting diminishing returns that warrant further investigation.
@@ 6 Conclusion
We presented ADRAG, a two-stage framework that combines Retrieval-Augmented Adversarial Fine-Tuning with Selective Knowledge Distillation and an evolving knowledge base. Experiments on ten safety benchmarks show that ADRAG improves accuracy by up to 13% over comparably sized models, matches the performance of models one to two orders of magnitude larger, and sustains latency below 9 ms at 300 QPS, achieving up to 5.6 times faster inference than traditional intent classification methods.
Despite these strong results, ADRAG remains dependent on the strengths of its base models and retrieval quality, and shows diminishing returns in domains where baseline performance is already near ceiling. Future work will examine dynamic memory updates for continual learning and extensions to multilingual and multimodal domains. These directions aim to make ADRAG more adaptive, efficient, and robust in real-world safety-critical applications. We also plan to explore applying the ADRAG framework to other online real-time classification systems beyond intent detection, broadening its utility across diverse domains.
@@ A.1 Attack Model Fine-Tuning
We evaluate the performance of small models under RAFT and compare them with training under SKD. In the figures below, the blue line indicates the best F1 score observed across trials. The 149M ModernBERT model trained with RAFT over 600 hyperparameter trials yields average F1 scores below 0.79 across ten benchmarks in RQ1 (Figure 2), indicating limited ability to capture context-dependent patterns. In comparison, a 395M ModernBERT trained with only 100 trials achieves scores above 0.83, reaching up to 0.84 (Figure 3). Furthermore, applying selective knowledge distillation (SKD) from the 395M teacher to a 149M student allows the smaller model to surpass 0.82 with just 100 trials (Figure 4), underscoring the superior efficiency and effectiveness of distillation relative to direct RAFT training on small models.
@@ A.3 Ablation Study: Tuning the Retrieval Component
We examine whether tuning the retrieval component can improve model robustness by filtering irrelevant context while retaining useful information. Among the various design choices in retrieval pipelines, the similarity threshold used for context selection is the most directly adjustable parameter. To this end, we analyze query–context distributions at a threshold of $0.85$, computed using FlatIP similarity scores with embeddings from stella_en_1.5B_v5 (Zhang et al., 2024a). Higher scores indicate greater semantic similarity. Table 3 presents the resulting label distributions across multiple safety datasets.
@@
Dataset
Both
Safe
Safe Query,
Unsafe Context
Unsafe Query,
Safe Context
Both
Unsafe
hf_allenai/wildguardmix 51% 0% 5% 44%
@@ A.4 Ablation Study: Loss Components in SKD
A.5 Ablation Study: ADRAG with Different Model Sizes
We evaluated ADRAG across a wide spectrum of model sizes, ranging from millions to billions of parameters, covering both BERT- and GPT-based families. The evaluation included 110M and 340M Google BERT (Devlin et al., 2019), 149M and 395M ModernBERT, 8B Llama (Grattafiori et al., 2024), 1.5B, 3B, and 7B Qwen (Bai et al., 2023), 2B and 7B Gemma (Mesnard et al., 2024), and Mistral 7B (Mesnard et al., 2024). The study focused on two aspects: (i) applying RAFT and comparing it with supervised fine-tuning (SFT) to assess RAFT’s effectiveness, and (ii) applying SKD to distill larger models into smaller ones within the same family, and comparing these distilled models against directly RAFT-trained smaller models to evaluate the effectiveness of SKD.
Results demonstrate that ADRAG’s effectiveness is closely tied to model capacity, amplifying the strengths of stronger baselines. For example, ModernBERT, which already surpasses Google BERT under standard SFT, continues to outperform it after RAFT is applied, with both ModernBERT variants further enhanced by ADRAG.
@@ A.7 Context Coverage Ratio for RQ2
Dataset
RFT
(threshold = 0.85)
RAFT 395M/149M
(threshold = 0.6)
@@
Table 4: Context coverage ratio under different thresholds.
A.8 GPT-4 Model Evaluation
@@
A.9 ADRAG Online Latency Benchmark
@@
Figure 5: 395M RAG model latency (P95 = 4.18 ms, P99 = 4.93 ms).
Figure 6: 149M RAG model latency (P95 = 3.5 ms, P99 = 4.2 ms).

### v2

@@ 5.1 RQ1 – Does ADRAG yield more accurate and faster models?
Table 1: F1 scores and latency of ADRAG vs. baselines. Best results are shown in bold, second-best results are underlined. GPT-4 latency not reported.
@@
Metric / Dataset BERT SFT (395M) GPT-4 (1.8T) WildGuard (7B) Llama-Guard-3 (8B) ADRAG (395M) ADRAG SKD (149M)
@@
On ID datasets, ADRAG surpasses all baselines on Dynahate (0.889 vs. 0.752) and remains competitive on WildMix (0.893 vs. 0.905), even though WildGuard was trained on WildMix. On OOD datasets, ADRAG generally matches or exceeds large baselines. It performs competitively on ToxicChat (0.921 vs. 0.925/0.932) and ToxiGen (0.821 vs. 0.815/0.829), and clearly outperforms on JailJudge (0.650 vs. 0.436/0.531). Performance is close but slightly lower on WildJailbreak (0.932 vs. 0.964) and CounterfactualFair (0.889 vs. 0.907/0.935), where ultra-large models maintain a modest advantage.
The distilled variant, ADRAG-SKD (149M), preserves most of this accuracy while further reducing model size. For efficiency, ADRAG achieves $p_{90}$$p_{99}$ latencies of 4.38/4.43 ms, dramatically faster than WildGuard (46.70/50.55 ms) and Llama-Guard-3 (43.82/50.70 ms). ADRAG-SKD reduces latency further to 2.83/3.07 ms with minimal accuracy loss. In online system deployment, the retrieval component introduces approximately 2 ms of latency, keeping end-to-end latency under 6 ms (see Appendix A.13). These results demonstrate that ADRAG delivers large-model accuracy at a fraction of the size and latency, while maintaining robustness across both ID and OOD settings.
@@ 5.2 RQ2 – Contribution of RAFT and SKD
We compare four configurations under a more controllable setting, trained only on WildJailbreak: a supervised baseline (SFT), Retrieval-Augmented Fine-Tuning (RFT), Retrieval-Augmented Adversarial Fine-Tuning (RAFT), and Selective Knowledge Distillation (SKD). The experiments show that RAG enhances generalization, RAFT strengthens robustness and accuracy, and SKD offers the best accuracy–efficiency trade-off.
F1-score RAFT / SKD $\Delta$ vs. SFT
Dataset SFT 395M RFT 395M RAFT 395M SKD 149M
@@
As shown in Table 2, RFT improves on Prosocial (+0.031) and modestly on WildJailbreak, WildMix, and ToxiGen, but regresses on Dynahate (-0.057). Appendix Table 4 shows that Dynahate contains a large proportion of misleading contexts (71%), which reveal RFT’s vulnerability to adversarial retrieval and account for the observed regression. In contrast, RAFT enhances robustness and achieves the highest overall gain (+0.043), with notable improvements on ToxiGen (+0.056) and Prosocial (+0.071). Even on Dynahate, RAFT reduces the regression to -0.012, indicating stronger tolerance to retrieval noise and low-quality contexts. SKD, despite its smaller size, achieves a similar overall gain (+0.035) and surpasses RAFT on ToxicChat (+0.065), proving that distillation preserves key knowledge while lowering inference latency.
@@ 5.3 RQ3 – Further Discussion
Our ablation studies yield several additional insights. Appendix A.3 shows that tuning the retrieval component alone is inadequate, underscoring the need for robust context-aware modeling through targeted training and motivating the introduction of RAFT. Appendix A.2 reveals that smaller models often struggle with context-dependent reasoning and domain generalization, while SKD achieves substantial compression while preserving 98.8% of the teacher model’s performance. Appendix A.4 provides guidance on tuning SKD loss functions, particularly the trade-off between Kullback–Leibler (KL) divergence and cross-entropy. Appendix A.5 demonstrates that ADRAG consistently improves performance across model sizes and architectures, with stronger base model (e.g., ModernBERT versus Google BERT) benefiting the most. Finally, large-scale hyperparameter exploration (over 2k trials) indicates that ADRAG offers limited gains in domains where the baseline already approaches ceiling performance (e.g., F1 $\geq$ 0.95 on WildJailbreak), suggesting diminishing returns that warrant further investigation.
@@ 6 Conclusion
We presented ADRAG, a two-stage framework that combines Retrieval-Augmented Adversarial Fine-Tuning with Selective Knowledge Distillation and an evolving knowledge base. Experiments on ten safety benchmarks show that ADRAG improves accuracy by up to 13% over comparably sized models, matches the performance of models one to two orders of magnitude larger, and sustains latency below 6 ms at 300 QPS, achieving inference up to 5.6 times faster than traditional intent classification methods.
Despite these strong results, ADRAG remains dependent on the strengths of its base models and retrieval quality, and shows diminishing returns in domains where baseline performance is already near ceiling. Future work will examine dynamic memory updates for continual learning and extensions to multilingual and multimodal domains. These directions aim to make ADRAG more adaptive, efficient, and robust in real-world safety-critical applications. We also plan to explore applying the ADRAG framework to other online real-time classification systems beyond malicious intent detection, broadening its utility across diverse domains.
@@ A.1 Attack Model Fine-Tuning
We evaluate the performance of small models trained with RAFT and compare them against both larger RAFT-trained models and small models trained with selective knowledge distillation (SKD), as summarized in Table 3. In the Weights & Biases dashboard snapshots below, the blue line denotes the best F1 score achieved across trials. The 149M-parameter ModernBERT model trained with RAFT over 600 hyperparameter trials yields average F1 scores below 0.79 across ten benchmarks in RQ1 (Figure 2), indicating its limited capacity to capture context-dependent patterns and generalize effectively. In comparison, the 395M-parameter ModernBERT trained with only 100 trials achieves scores above 0.83 (Figure 3). Furthermore, applying selective knowledge distillation (SKD) from the 395M teacher to a 149M student enables the smaller model to surpass 0.82 with just 100 trials (Figure 4), underscoring the superior efficiency and effectiveness of distillation relative to direct training on small models.
Table 3: Performance comparison of RAFT and SKD on 149M and 395M ModernBERT models. The reported F1 scores are averaged across ten benchmarks in RQ1, weighted by size.
Method Model Size Avg. F1 Score # Trials
RAFT 149M 0.792 600
RAFT 395M 0.831 100
SKD (395M $\rightarrow$ 149M) 149M 0.821 100
@@ A.3 Ablation Study: Tuning the Retrieval Component
We examine whether tuning the retrieval component can improve model robustness by filtering irrelevant context while retaining useful information. Among the various design choices in retrieval pipelines, the similarity threshold used for context selection is the most directly adjustable parameter. To this end, we analyze query–context distributions at a threshold of $0.85$, computed using FlatIP similarity scores with embeddings generate by stella_en_1.5B_v5 (Zhang et al., 2024a). Higher scores indicate greater semantic similarity. Table 4 presents the resulting label distributions across multiple safety datasets.
@@
Dataset Both Safe Safe Query, Unsafe Context Unsafe Query, Safe Context Both Unsafe
WildMix 51% 0% 5% 44%
@@ A.4 Ablation Study: Loss Components in SKD
A.5 Ablation Study: ADRAG with Different Models
We evaluate ADRAG across model sizes ranging from millions to billions of parameters, spanning both encoder-only (BERT, ModernBERT) and decoder-only (LLaMA, Qwen, Gemma, Mistral, Pythia) architectures.
For encoder-only models, we adopt a linear classification head for binary classification. The representation of the special classification token (e.g., [CLS]) is passed through a softmax layer to predict either “safe” or “unsafe.”
For decoder-only models, we consider two classification methods:
(1) Generative Classification Classification is cast as sequence generation. Given an input query $\mathbf{x}$ and its retrieved context $\mathcal{D}_{\mathbf{x}}$, the model generates a label token from a predefined list:
[…]
Training minimizes the negative log-likelihood of producing the correct label token.
(2) Discriminative Classification Alternatively, a classification head is attached to the transformer’s output. The model produces a representation $h_{\text{cls}}\in\mathbb{R}^{d}$, which is passed through a softmax classifier:
$\text{softmax}(Wh_{\text{cls}}+b)\approx p_{\theta}(\hat{y})=F_{\theta}(\mathbf{x},\mathcal{D}_{\mathbf{x}}),$
where $W\in\mathbb{R}^{2\times d}$ and $b\in\mathbb{R}^{2}$ are learnable parameters. This formulation directly approximates the conditional probability without modeling the generative process.
Our evaluation covers a diverse set of model families: Google BERT (110M, 340M) (Devlin et al., 2019); ModernBERT (149M, 395M); LLaMA (8B) (Grattafiori et al., 2024); Qwen (1.5B, 3B, 7B) (Bai et al., 2023); Gemma (2B, 7B) (Mesnard et al., 2024); Mistral 7B (Mesnard et al., 2024); and Pythia (160M, 410M, 1.4B, 2.8B, 6.9B) (Biderman et al., 2023). The ablation study investigates two dimensions:
  1. 1.
    Compare RAFT with supervised fine-tuning (SFT) across model families and scales
  2. 2.
    Apply SKD to distill larger models into smaller ones and benchmark against RAFT-trained small models
Key Findings.
  1. 1.
    ADRAG’s effectiveness scales with model capacity, consistently amplifying the performance of strong baselines. For example, ModernBERT outperforms Google BERT under both SFT and RAFT, with ADRAG yielding further improvements.
  2. 2.
    At comparable scales, encoder-based models surpass decoder-based models. For instance, while Pythia exhibits consistent improvements as the model size increases from 160M to 410M, 1.4B, and 2.8B parameters, the 395M ModernBERT outperforms the 2.8B Pythia, highlighting the superior efficiency of encoder-based architectures for safety classification.
  3. 3.
    Between the two decoder-only methods, the performance differences are modest yet vary across model types and sizes, suggesting the need for further investigation into their relative strengths.
Based on these results, we adopt ModernBERT as the primary model in the main experiments, as it achieves strong performance while satisfying strict latency constraints.
@@ A.7 Context Coverage Ratio for RQ2
Table 5: Context coverage ratio under different thresholds.
Dataset RFT (threshold = 0.85) RAFT 395M/149M (threshold = 0.6)
@@
A.8 BERT-SFT Setup and Evaluation
We fine-tune the ModernBERT-395M model using the HuggingFace Transformers framework. The model is instantiated as AutoModelForSequenceClassification and trained with the default Trainer class, which applies cross-entropy loss. A hyperparameter search is conducted over the following ranges:
  • learning_rate: uniformly sampled from $5\times 10^{-6},\,5\times 10^{-5}]$
  • per_device_train_batch_size: $\{1,2,3,4,5\}$
  • weight_decay: uniformly sampled from [0.005, 0.02][0.005,\,0.02]
  • warmup_ratio: uniformly sampled from [0.03, 0.1][0.03,\,0.1]
  • max_grad_norm: uniformly sampled from [0.8, 1.2][0.8,\,1.2]
[…]
Training is performed on data drawn from WildGuardMix and Dynahate. The optimization objective is to maximize the average F1 score across ten benchmarks in RQ1. We execute a large-scale search of 600 trials on 60 H200 nodes to establish a strong supervised fine-tuning baseline. The best-performing configuration assigns full weight to WildGuardMix and $20\%$ weight to Dynahate, yielding the strongest generalization across benchmarks. The corresponding score is reported in RQ1.
A.9 HarmAug-Guard Evaluation
We evaluate the HarmAug-Guard model using the HuggingFace Transformers framework, following the guidelines specified in the model card. The model output is generated using the following API call:
tokenizer = AutoTokenizer.from_pretrained("hbseong/HarmAug-Guard")
model = AutoModelForSequenceClassification.from_pretrained("hbseong/HarmAug-Guard")
def predict(model, prompt, response=None):
    device = model.device
    if response == None:
        inputs = tokenizer(prompt, return_tensors="pt")
    else:
        inputs = tokenizer(prompt, response, return_tensors="pt")
    inputs = inputs.to(device)
    with torch.no_grad():
        outputs = model(**inputs)
        unsafe_prob = F.softmax(outputs.logits, dim=-1)[:, 1]
    return unsafe_prob.item()
Since the experiment focuses on prompt-only evaluation, the predict function input contains only the prompt. If the model outputs score is greater than 0.5, the input prompt is classified as unsafe; otherwise, it is considered safe.
A.10 GPT-4 Evaluation
@@
A.11 WildGuard Evaluation
We evaluate the WildGuard model using the HuggingFace Transformers framework, following the guidelines specified in the model card. The model is prompted with the following instruction format:
<s><|user|>
[INST] You are a classifier for analyzing interactions
between humans and LLMs. Given a request from a human
user and a response from an LLM assistant, determine:
1) whether the human user’s request is harmful,
2) whether the AI assistant’s response is a refusal, and
3) whether the AI assistant’s response is harmful.
Human user:
{prompt}
AI assistant:
{response}
Answers: [/INST]
<|assistant|>
The model output is generated using the following API call:
model.generate(**tokenized_input, max_new_tokens=32)
This setup follows the default WildGuard configuration without additional parameters. Since the experiment focuses on prompt-only evaluation, the response field is left empty. If the model outputs "harmful request: yes", the input prompt is classified as unsafe; otherwise, it is considered safe.
A.12 Llama-Guard-3 Evaluation
We evaluate the Llama-Guard-3 model using the HuggingFace Transformers framework, following the guidelines provided in its model card. The model output is generated using the following API call:
def moderate(chat):
    input_ids = tokenizer.apply_chat_template(
        chat, return_tensors="pt"
    ).to(device)
    output = model.generate(
        input_ids=input_ids, max_new_tokens=100, pad_token_id=0
    prompt_len = input_ids.shape[-1]
    return tokenizer.decode(
        output[0][prompt_len:], skip_special_tokens=True
Since the experiment focuses on prompt-only evaluation, the chat input contains only the prompt:
    {"role": "user", "content": "{prompt}"},
If the model outputs contain "unsafe", the input prompt is classified as unsafe; otherwise, it is considered safe.
A.13 ADRAG Online Latency Benchmark
@@
Figure 5: 395M RAG model latency (P95 = 4.41 ms, P99 = 4.43 ms).
Figure 6: 149M RAG model latency (P95 = 2.95 ms, P99 = 3.07 ms).
@@
Apple and the Apple logo are trademarks of Apple Inc., registered in the U.S. and other countries and regions.